Fermi National Laboratory


Accelerator Update for February 2, 2005 - February 4, 2005

Wednesday February 2
The day shift began with Operations monitoring store 3956, with the Antiproton Source (Pbar) stacking, and with MiniBooNE and the Meson experiment MCenter taking beam.

At 8:19 AM, quadrupole power supply (I:Q703) tripped off. This put a halt to stacking. Experts began their repairs.

At 9:48 AM, a Tevatron (TeV) RF station (TRF2) tripped off on Low DC Voltage. It was left off until after the store ends.

Power supply experts reported at 12:16 PM that they had replaced the Q703 power supply.

Pbar resumed stacking at 12:35 PM.

Experimenters at MTest asked for beam at 12:49 PM.

NuMI experts began a study around 5:35 PM.

A Pbar expert began a stacking study at 6 PM.

MTest experimenters reported at 10:18 PM that they were finished with the beam for the night.

At 11:18 PM, I:Q703 tripped off, but it reset okay.

Thursday February 3
The midnight shift began with Operations monitoring store 3956, with Pbar stacking, and with MiniBooNE and MCenter taking beam.

At 5:50 AM, I:Q703 tripped off. It reset okay.

Operations terminated store 3956 at 8 AM.

At 8:17 AM, I:Q703 tripped off and wouldn't reset. Operators notified an expert.

Cryo system technicians began replacing the TeV sector C3 wet engine flywheel at 8:46 AM.

Power supply techs completed their repair of I:Q703 at 9:55 PM. Stacking resumed, as did Switchyard (SY) beam.

Cryo techs reported at 10:19 AM that they had completed the C3 wet engine work.

An RF expert reported at 10:46 AM that TRF2 was back up and running.

Operations put the TeV into shot setup at 11:36 AM.

Operations established store 3958 at 2:22 PM with an initial luminosity of 104.0E30. Pbar resumed stacking.

Operations resumed sending beam to MCenter and MTest at 2:54 PM.

At 3:52 PM, the TeV suffered a sector E1 quench. Experts investigated and found a problem the T:VA35 regulator.

NuMI experts began taking beam at 4:25 PM and sending it to the Main Injector (MI) dump.

Operations made the Collider Detector at Fermilab (CDF) and D-Zero (D0) ready for access around 6 PM. They were both out of access about two hours later.

NuMI experts finished their studies at 10 PM.

TeV experts began a study period at 11:49 PM.

Friday February 4
The midnight shift began with TeV experts conducting octupole studies, with Pbar stacking, and with MiniBooNE taking beam.

At 12:40 AM, the TeV suffered a sector DQ0 quench.

Operations put the TeV into shot setup at 5:48 AM.

The Plan for Friday and the Weekend
The plan for today and this weekend is to stack, stash, and store.


Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.



last modified 2/4/2005   email Fermilab